# Periodic classical trajectories and quantum scars in many-spin systems

### v2

@@ V Classical spins: transient nearly quasiperiodic regime
While computing the Lyapunov exponents for periodic trajectories, we discovered that, as the Lyapunov instability develops and the deviations from the reference periodic trajectory become large and no longer describable by Lyapunov exponents, the dynamics does not immediately become ergodic — rather it often enters a nearly quasiperiodic regime illustrated in Fig. 9. The lifetime of this regime exhibits strong variations: it may be relatively short — a few time units (one unit is $11/h=$ 1), or fairly long — a few hundred time units — as is the case in Fig. 9(b), or anomalously long — meaning $\sim 10^{6}$ time units or more; one such example is shown in Figs. 9(a) and 10. Later in this section we identify the anomalously long-lived quasiperiodic regime with the phenomenon of Arnold diffusion. In our simulations we came across only five cases of this regime, which are marked as such in Figs.4(b,c).
@@ V.1 Frequency spectra
Given the discussion in the subsection IV.5.2 of the translational symmetry breaking accompanying Lyapunov instabilities, the explanation of the origin of the nearly quasiperiodic regime can be the following. The initial periodic motion corresponds to the wave number $q=0$. The motion that emerges as a result of the Lyapunov instability is characterised by wave vector $q=q_{\text{p}}\neq 0$. As the Lyapunov instability develops, the spectral weight becomes gradually transferred from the component with the wave number $q=0$ to the one with $q=q_{\text{p}}$. (The latter mode can be treated as a superposition of two complex-valued waves proportional to $e^{\pm iq_{\text{p}}r}$ and forming a real-valued standing wave.) The growth of the second mode saturates when the amplitude of that mode becomes comparable with that of the $q=0$ mode, while the modes at other $q$ are still too small. The phase space of thus emerging dynamics becomes effectively four-dimensional. Systems with four-dimensional phase spaces often have finite regions of non-chaotic quasiperiodic dynamics separated from the regions of stochastic dynamics[126, 43, 127]. This is because a quasiperiodic trajectory in a four-dimensional phase space, when present, covers a two-dimensional torus within a three-dimensional energy shell, thereby dividing the energy shell into two parts with the border impenetrable for stochastic trajectories. When a trajectory ends up in a stochastic regime, it can be strongly or weakly chaotic. A weakly chaotic trajectory spends a long time in the vicinity of a truly quasiperiodic trajectory. We believe that our numerical evidence of the quasiperiodic behavior is of the latter kind.
@@
The motion departing from an unstable periodic trajectory is expected to end up in the stochastic part in the above 4-dimensional phase space on the basis of the following argument. Consider a two-dimensional Poincare surface of section (PSS) perpendicular to the unstable periodic trajectory within the three-dimensional energy shell. The point where that trajectory crosses the PSS is an unstable fixed point for the corresponding 2D dynamical map. An unstable fixed point in a 2D subspace should be accompanied by a separatrix. The motion in a vicinity of a separatrix is supposed to form a stochastic layer, because the time between the subsequent crossings of the PSS diverges on the approach to the separatrix, implying that there is an increasingly dense set of orbits satisfying the resonance condition $T_{1}=nT_{0}$, where $T_{0}$ and $T_{1}$ are, respectively, the shorter and the longer periods on the KAM-torus and $n$ is an integer. As a result, a stochastic layer emerges, and our trajectory should find itself “stuck” within that layer. Such a regime would mean a suppressed ergodicity with a strong quasiperiodic tendency. In particular, the the frequency of the slower periodic motion would exhibit a drift over time that is longer than the already long period of that motion. This regime would persist until the four-dimensional character of the dynamics is destroyed by the mechanisms to be introduced in the next part.
@@ V.4 Time evolution of Fourier modes with different wave numbers $q$
We further note that the exponential onset of each $q=q_{p}$ mode in Fig. 12 is accompanied by the subsequent onset of the modes growing as $F_{q}(t)\propto\exp({2n\lambda_{\text{p}}t})$, where $n$ is a natural number. These modes are explicitly indicated by the respective exponential fits in Fig. 12(c). The straight lines representing these exponential fits on a semilog plot tend to intersect roughly at the same point located on the red line representing the mode $F_{0}(t)$. This phenomenology can be explained as follows. The growth of the leading $q=q_{p}$ mode induces dominant modulation of the chain background. As a result, the spatial Fourier harmonics with $q=nq_{p}$ become coupled to the $q=q_{p}$ mode in the $n$th order of the perturbation theory, and their amplitudes emerge at later times out of statistical noise exhibiting the Lyapunov growth with the rate $n\lambda_{p}$, [$2n\lambda_{p}$ for $F_{q}(t)$]. This phenomenology, however, does not signify independent modes – rather they are perturbation-induced admixtures to the original unstable $q=q_{p}$ mode. That perturbation can be described in terms of backfolding of the chain’s Brillouin zone, which is to be discussed further below in subsection V.6 in the context of Mechanism B.
@@
Figure 13: Illustration of criterion for the onset of Mechanism A that destroys the nearly quasiperiodic regime for spin chains with $J=1.76$ and various lengths $L$. The upper plot $\lambda(q)$ shows parabolic approximations (19) and (20) for two ranges of wave numbers $q$ of the Fourier modes 18 possessing positive Lyapunov exponents $\lambda$ that lead to instabilities around the original periodic trajectory. The parameters for the larger range (19) are given in Fig. 4(b). The smaller range (20) is obtained with $q_{1}=1.51$, $\lambda_{\text{1,max}}=0.104$ and $\alpha_{1}=1.23\cdot 10^{2}$ extracted from the fit to the numerical data shown in the inset. The lengths $L=6,18,19,21,42$ correspond to chains whose Fourier modes are plotted in Fig. 12. The horisontal lines next to each $L$ represent the $q$-axes with allowed values $q=\frac{2\pi}{L}n$ marked and the unstable ranges magenta-colored. In each case, there is, at least, one allowed $q$-value falling within an unstable range and responsible for the primary Lyapunov instability around the original periodic trajectory and the subsequent onset of the nearly quasiperiodic regime. The onset of Mechanism A requires at least one more allowed $q$ within an unstable range. Such $q$ is present for $L=21$ and $L=42$, hence Mechanism A is operational — see Figs. 12(a1,a2). For $L=6,18$ and 19, there are no additional values $q=\frac{2\pi}{L}n$ in the unstable ranges, hence the Mechanism A is not operational — see Figs. 12(b1,b2,c).
@@ V.7 Mechanism C: Arnold diffusion
For the initial condition “all spins up”, Eq.(2), the integrable Hamiltonian (22) generates the same periodic trajectory for each spin as our original Hamiltonian (III). We now reexpress the latter as
@@
The term ${\cal H}_{1}$ is, in general, not a small perturbation to ${\cal H}_{0}$. However, in the vicinity of the considered periodic trajectory it is.
@@
The transient character of this regime implies that the strict application of the limit $t\to\infty$ to the definition (9) of Lyapunov exponent implies that that definition would give the same $\lambda_{\text{max}}$ in both Arnold diffusion and the ergodic regime.
@@ VI Quantum spins
As discussed in the introduction, the present work was initially motivated by our interest to explore whether periodic orbits in classical spin lattices imply the existence of many-body scar eigenstates in quantum spin lattices. Our classical investigation in Section IV indicates that the chances of observing the scar effects in many-spin systems are, actually, somewhat better than what one could expect from the knowledge of the Lyapunov spectra of ergodic trajectories[120]. The reason is that one can choose the parameters of the Hamiltonian such that the largest Lyapunov exponent for a periodic trajectory is significantly smaller than that for an ergodic trajectory (cf. Figs. 3 and 23), and, moreover, there exist rather long spin chains [see Figs. 4(b,c)] for which the Lyapunov instabilities in the vicinity of the periodic trajectories are completely suppressed.
Below, we explore the quantum dynamics starting with the special initial wave function $|\Psi^{\text{up}}\rangle$ –“all spins up” defined by Eq.(3), which in the classical limit gives a periodic trajectory. We first check whether the evolution starting from $|\Psi^{\text{up}}\rangle$ exhibits an anomalously slow thermalisation. Then we decompose $|\Psi^{\text{up}}\rangle$ into the energy eigenstates and check whether any of the eigenstates prominently present in the above initial state also violates eigenstate thermalisation hypothesis.
@@ VI.1 Suppressed initial thermalization
Figure 16: Slowdown of thermalisation dynamics illustrated by the relaxation of $\langle S^{z}_{1}(t)\rangle$ for chains of quantum spins $S$ with lengths $L$; $\tilde{J}=1.76$. Blue solid lines and orange dashed lines represent quantum relaxation starting, respectively, from the initial states $|\Psi^{\text{up}}\rangle$ and $|\Psi^{\text{inf}}\rangle$. In all cases, the generic relaxation curves starting from $|\Psi^{\text{inf}}\rangle$ quickly approach zero, exhibiting afterwards only small fluctuations. In contrast, the relaxation curves starting from $|\Psi^{\text{up}}\rangle$ exhibit slow decaying positively-valued long-time tails for chains with $S\geq 1$. (To improve computational efficiency, the relaxation curves for $|\Psi^{\text{inf}}\rangle$ were computed for smaller chain lengths than indicated in the legends, namely, $L=$ 13, 8, 7 and 6 for $S=$ 1/2, 1, 3/2 and 2, respectively. This affected only small fluctuations of $\langle S^{z}_{1}(t)\rangle$ around zero.) Brown dash-dotted lines represent classical simulations averaged over ensembles of $10^{4}$ initial conditions imitating the quantum state $|\Psi^{\text{up}}\rangle$ as described in the text.
Figure 17: Slowdown of thermalisation dynamics for chains of quantum spins 3/2 with different lengths $L$ for the initial state $|\Psi^{\text{up}}\rangle$. Plotted is the relaxation of $\langle S^{z}_{1}(t)\rangle$ with $\tilde{J}=1.76$, same as in Fig. 16. The closeness of the relaxation curves for different $L$ is indicative of a the thermodynamic limit $L\to\infty$. Also included is the plot of generic relaxation starting with $|\Psi^{\text{inf}}\rangle$.
To characterise the initial thermalisation, we compute the quantum mechanical expectation value $\langle S^{z}_{1}(t)\rangle\equiv\langle\Psi(t)|S^{z}_{1}(t)|\Psi(t)\rangle$ starting from the fully polarized state $|\Psi^{\text{up}}\rangle$ given by Eq.(3) and compare it with the thermalisation of the same observable starting from a more generic initial condition where the spin $S^{1}_{z}$ is fully polarised but the rest of the system is at the infinite temperature equilibrium:
@@
where $|\text{inf}\rangle_{L-1}$ is a pure state randomly sampled in the Hilbert space of spins on sites from 2 to $L$  [^1] A pure quantum state randomly sampling the infinite temperature equilibrium in a Hilbert space of dimension $N\gg 1$ is obtained as $|\text{inf}\rangle=\frac{1}{\sqrt{2^{L-1}}}\sum^{2^{L-1}}_{n=0}c_{n}|b_{n}\rangle$, where $|b_{n}\rangle$ are basis states, and $c_{n}=|c_{n}|e^{i\phi_{n}}$ are complex amplitudes with phases $\phi_{n}$ chosen randomly from the interval $[0,2\pi)$ and absolute values $|c_{n}|$ chosen using the probability distribution $P(|c_{n}|^{2})=Ne^{-N|c_{n}|^{2}}$ (see Refs.[162, 163]). Both $|\Psi^{\text{up}}\rangle$ and $|\Psi^{\text{inf}}\rangle$ have the same energy expectation values, thereby representing the same “energy shell” in the sense of a canonical ensemble.
@@
The relaxation plots of $\langle S^{z}_{1}(t)\rangle$ for $S=$ 1/2, 1, 3/2 and 2 are shown in Fig. 16. There one can observe that the thermalisation process for $S\geq 3/2$ starting from the initial condition $|\Psi^{\text{up}}\rangle$ is noticeably slower than the one starting from $|\Psi^{\text{inf}}\rangle$. Such a difference is to be expected when atypical quantum scar eigenstates are prominently present in the expansion of $|\Psi^{\text{up}}\rangle$. At the same time, the difference between the two thermalisation curves for $S=1/2$ is rather small — suggestive of the absence of quantum scar eigenstates in the expansion of $|\Psi^{\text{up}}\rangle$. The case of $S=1$ appears to be transitional between $S=1/2$ and $S\geq 3/2$.
@@
Figure 17 further illustrates the anomalous slowdown of the thermalisation for spin-3/2 chains as a function of chain length $L$. The coincidence of the initial slowed down behavior for the chains of different length indicates that the slowdown of the thermalisation process starting from the state $|\Psi^{\text{up}}\rangle$ is a feature that is also present in the thermodynamic limit $L\to\infty$.
@@
We, finally, note that Fig. 16 also includes the results of classical simulations for an ensemble of initial conditions devised to imitate the relaxation of quantum spins $S$ starting from the initial state $|\Psi^{\text{up}}\rangle$. In the simulations, each classical spin has length 1. In order to represent a quantum state $|m=S\rangle$, the range $[-1,1]$ of the $z$-projection of a classical spin is divided into $12S+$ 1 equal intervals of length $\frac{2}{2s+1}$ and then the classical projection $S_{z}$ is randomly sampled from the topmost interval $[1-\frac{2}{2s+1},1]$, while the $x$- and $y$-projections are chosen as $S_{x}=\sqrt{1-S_{z}^{2}}\cos\varphi$ and $S_{y}=\sqrt{1-S_{z}^{2}}\sin\varphi$, where $\varphi$ is a random angle from the interval $[0,2\pi)$. On the one hand, the above ensemble becomes increasingly close to the considered periodic classical trajectories as the value of quantum spin $S$ increases. On the other hand, as one can see in Fig. 16, the average for that ensemble exhibits an excellent agreement with the quantum simulations starting already with $S=1$. This is yet another indication, in addition to the observations of Refs.[68], that classical spin ensembles can quantitatively represent quantum spin dynamics for translationally invariant chains of spins $S\geq 1$ at high temperatures.

### v3

@@ V Classical spins: transient nearly quasiperiodic regime
While computing the Lyapunov exponents for periodic trajectories, we discovered that, as the Lyapunov instability develops, the deviations from the reference periodic trajectory become large and no longer describable by Lyapunov exponents, and yet the dynamics does not immediately become ergodic — rather it often enters a nearly quasiperiodic regime illustrated in Fig. 9. The lifetime of this regime exhibits strong variations: it may be relatively short — a few time units (one unit is $11/h=$ 1), or fairly long — a few hundred time units — as is the case in Fig. 9(b), or anomalously long — meaning $\sim 10^{6}$ time units or more; one such example is shown in Figs. 9(a) and 10. Later in this section we identify the anomalously long-lived quasiperiodic regime with the phenomenon of Arnold diffusion. In our simulations we came across only five cases of this regime, which are marked as such in Figs.4(b,c).
@@ V.1 Frequency spectra
Given the discussion in the subsection IV.5.2 of the translational symmetry breaking accompanying Lyapunov instabilities, the explanation of the origin of the nearly quasiperiodic regime can be the following. The initial periodic motion corresponds to the wave number $q=0$. The motion that emerges as a result of the Lyapunov instability is characterised by wave vector $q=q_{\text{p}}\neq 0$. As the Lyapunov instability develops, the spectral weight becomes gradually transferred from the component with the wave number $q=0$ to the one with $q=q_{\text{p}}$. (The latter mode can be treated as a superposition of two complex-valued waves proportional to $e^{\pm iq_{\text{p}}r}$ and forming a real-valued standing wave.) The growth of the second mode saturates when the amplitude of that mode becomes comparable with that of the $q=0$ mode, while the modes at other $q$ are still too small. The phase space of the emerging dynamics becomes effectively four-dimensional. Systems with four-dimensional phase spaces often have finite regions of non-chaotic quasiperiodic dynamics separated from the regions of stochastic dynamics[136, 43, 137]. This is because a quasiperiodic trajectory in a four-dimensional phase space, when present, covers a two-dimensional torus within a three-dimensional energy shell, thereby dividing the energy shell into two parts with the border impenetrable for stochastic trajectories. When a trajectory ends up in a stochastic regime, it can be strongly or weakly chaotic. A weakly chaotic trajectory spends a long time in the vicinity of a truly quasiperiodic trajectory. We believe that our numerical evidence of the quasiperiodic behavior is of the latter kind.
@@
The motion departing from an unstable periodic trajectory is expected to end up in the stochastic part in the above 4-dimensional phase space on the basis of the following argument. Consider a two-dimensional Poincaré surface of section (PSS) perpendicular to the unstable periodic trajectory within the three-dimensional energy shell. The point where that trajectory crosses the PSS is an unstable fixed point for the corresponding 2D dynamical map. An unstable fixed point in a 2D subspace should be accompanied by a separatrix. The motion in a vicinity of a separatrix is supposed to form a stochastic layer[43], because the time between the subsequent crossings of the PSS diverges on the approach to the separatrix, implying that there is an increasingly dense set of orbits satisfying the resonance condition $\mathcal{T}_{1}=n\mathcal{T}_{0}$, where $\mathcal{T}_{0}$ and $\mathcal{T}_{1}$ are, respectively, the shorter and the longer periods on the KAM-torus and $n$ is an integer. As a result, a stochastic layer emerges, and our trajectory should find itself “stuck” within that layer. Such a regime would mean a suppressed ergodicity with a strong quasiperiodic tendency. In particular, the frequency of the slower periodic motion would exhibit a drift over time that is longer than the already long period of that motion. This regime would persist until the four-dimensional character of the dynamics is destroyed by the mechanisms to be introduced in the next part.
@@ V.4 Time evolution of Fourier modes with different wave numbers $q$
We further note that the exponential onset of each $q=q_{p}$ mode in Fig. 12 is accompanied by the subsequent onset of the modes growing as $F_{q}(t)\propto\exp({2n\lambda_{\text{p}}t})$, where $n$ is a natural number. These modes are explicitly indicated by the respective exponential fits in Fig. 12(c). The straight lines representing these exponential fits on a semilog plot tend to intersect roughly at the same point located on the red line representing the mode $F_{0}(t)$. This phenomenology can be explained as follows. The growth of the leading $q=q_{p}$ mode induces a modulation of the chain background, which, on the one hand, constitutes a dominant perturbation to the original translationally invariant periodic trajectory but, on the other hand, still remains very small as long as $F_{q_{p}}(t)\ll F_{0}(t)$. As a result, the spatial Fourier harmonics with $q=nq_{p}$ become coupled to the $q=q_{p}$ mode in the $n$th order of the perturbation theory, and their amplitudes emerge at later times out of statistical noise exhibiting the Lyapunov growth with the rate $n\lambda_{p}$, [$2n\lambda_{p}$ for $F_{q}(t)$]. This phenomenology, however, does not signify independent modes – rather they are perturbation-induced admixtures to the original unstable $q=q_{p}$ mode. That perturbation can be described in terms of backfolding of the chain’s Brillouin zone, which is to be discussed further below in subsection V.6 in the context of Mechanism B.
@@
Figure 13: Illustration of criterion for the onset of Mechanism A that destroys the nearly quasiperiodic regime for spin chains with $J=1.76$ and various lengths $L$. The upper plot $\lambda(q)$ shows parabolic approximations (20) and (21) for two ranges of wave numbers $q$ of the Fourier modes 19 possessing positive Lyapunov exponents $\lambda$ that lead to instabilities around the original periodic trajectory. The parameters for the larger range (20) are given in Fig. 4(b). The smaller range (21) is obtained with $q_{1}=1.51$, $\lambda_{\text{1,max}}=0.104$ and $\alpha_{1}=1.23\cdot 10^{2}$ extracted from the fit to the numerical data shown in the inset. The lengths $L=6,18,19,21,42$ correspond to chains whose Fourier modes are plotted in Fig. 12. The horisontal lines next to each $L$ represent the $q$-axes with allowed values $q=\frac{2\pi}{L}n$ marked and the unstable ranges magenta-colored. In each case, there is, at least, one allowed $q$-value falling within an unstable range and responsible for the primary Lyapunov instability around the original periodic trajectory and the subsequent onset of the nearly quasiperiodic regime. The onset of Mechanism A requires at least one more allowed $q$ within an unstable range. Such $q$ is present for $L=21$ and $L=42$, hence Mechanism A is operational — see Figs. 12(a1,a2). For $L=6,18$ and 19, there are no additional values $q=\frac{2\pi}{L}n$ in the unstable ranges, hence the Mechanism A is not operational — see Figs. 12(b1,b2,c).
@@ V.7 Mechanism C: Arnold diffusion
For the initial condition (3) “all spins up”, the integrable Hamiltonian (23) generates the same periodic trajectory for each spin as our original Hamiltonian (III). We now reexpress the latter as
@@
The term ${\cal H}_{1}$ is, in general, not a small perturbation to ${\cal H}_{0}$. However, in the vicinity of the considered periodic trajectory it is, because the right-hand-side of Eq.(V.7) becomes equal to zero once all spins pointing in the same direction and $J_{0}=2J/L$ are substituted into it.
@@
The transient character of this regime implies that the strict application of the limit $t\to\infty$ to the definition (10) of the Lyapunov exponent would give the same $\lambda_{\text{max}}$ in both Arnold diffusion and ergodic regime.
@@ VI Quantum spins
As discussed in the introduction, the present work was initially motivated by our interest to explore whether periodic orbits in classical spin lattices imply the existence of many-body scar eigenstates in quantum spin lattices. A quantum scar in a semiclassical limit can be associated with self-interference of a quantum wavepacket propagating along a periodic classical trajectory and thereby forming an atypical “standing wave”[115]. In a many-dimensional phase space, the presence of many Lyapunov instabilities along that trajectory is expected to make the above wavepacket too broad before it could start interfering with itself, thereby reducing the chances for a distinct scar eigenstate to emerge. From such a perspective, the weaker the Lyapunov instabilities, the better are the chances for observing quantum scars. In this regard, our classical investigation in Section IV indicates that the chances of observing the scar effects in many-spin systems are, actually, somewhat better than what one could expect from the knowledge of the Lyapunov spectra of ergodic trajectories[126]. The reason is that one can choose the parameters of the Hamiltonian such that the largest Lyapunov exponent for a periodic trajectory is significantly smaller than that for an ergodic trajectory (cf. Figs. 3 and 24), and, moreover, there exist rather long spin chains [see Figs. 4(b,c)] for which the Lyapunov instabilities in the vicinity of the periodic trajectories are completely suppressed.
Below, we explore the quantum dynamics starting with the special initial wave function $|\Psi^{\text{up}}\rangle$ –“all spins up” defined by Eq.(4), which in the classical limit gives a periodic trajectory. We first check whether the evolution starting from $|\Psi^{\text{up}}\rangle$ exhibits an anomalously slow thermalization. Then we decompose $|\Psi^{\text{up}}\rangle$ into the energy eigenstates and check whether any of the eigenstates prominently present in the above initial state also violates eigenstate thermalization hypothesis.
@@ VI.1 Suppressed initial thermalization
Refer to caption(a)(b)(c)(d)
Figure 16: Slowdown of thermalization dynamics illustrated by the relaxation of $\langle S^{z}_{1}(t)\rangle$ for chains of quantum spins $S$ with lengths $L$; $\tilde{J}=1.76$. Blue solid lines and orange dashed lines represent quantum relaxation starting, respectively, from the initial states $|\Psi^{\text{up}}\rangle$ and $|\Psi^{\text{inf}}\rangle$. In all cases, the generic relaxation curves starting from $|\Psi^{\text{inf}}\rangle$ quickly approach zero, exhibiting afterwards only small fluctuations. In contrast, the relaxation curves starting from $|\Psi^{\text{up}}\rangle$ exhibit slow decaying positively-valued long-time tails for chains with $S\geq 1$. (To improve computational efficiency, the relaxation curves for $|\Psi^{\text{inf}}\rangle$ were computed for smaller chain lengths than indicated in the legends, namely, $L=$ 13, 8, 7 and 6 for $S=$ 1/2, 1, 3/2 and 2, respectively. This affected only small fluctuations of $\langle S^{z}_{1}(t)\rangle$ around zero.) Brown dash-dotted lines represent classical simulations averaged over ensembles of $10^{4}$ initial conditions imitating the quantum state $|\Psi^{\text{up}}\rangle$ as described in the text.
Figure 17: Slowdown of thermalization dynamics for chains of quantum spins 3/2 with different lengths $L$ for the initial state $|\Psi^{\text{up}}\rangle$. Plotted is the relaxation of $\langle S^{z}_{1}(t)\rangle$ with $\tilde{J}=1.76$, same as in Fig. 16. The closeness of the relaxation curves for different $L$ is indicative of a the thermodynamic limit $L\to\infty$. Also included is the plot of generic relaxation starting with $|\Psi^{\text{inf}}\rangle$.
To characterise the initial thermalization, we compute the quantum mechanical expectation value $\langle S^{z}_{1}(t)\rangle\equiv\langle\Psi(t)|S^{z}_{1}(t)|\Psi(t)\rangle$ starting from the fully polarized state $|\Psi^{\text{up}}\rangle$ given by Eq.(4) and compare it with the thermalization of the same observable starting from a more generic initial condition where the spin $S^{1}_{z}$ is fully polarised but the rest of the system is at the infinite temperature equilibrium:
@@
where $|\text{inf}\rangle_{L-1}$ is a pure state randomly sampled in the Hilbert space of spins on sites from 2 to $L$  [^1] A pure quantum state randomly sampling the infinite temperature equilibrium in a Hilbert space of dimension $N\gg 1$ is obtained as $|\text{inf}\rangle=\frac{1}{\sqrt{2^{L-1}}}\sum^{2^{L-1}}_{n=0}c_{n}|b_{n}\rangle$, where $|b_{n}\rangle$ are basis states, and $c_{n}=|c_{n}|e^{i\phi_{n}}$ are complex amplitudes with phases $\phi_{n}$ chosen randomly from the interval $[0,2\pi)$ and absolute values $|c_{n}|$ chosen using the probability distribution $P(|c_{n}|^{2})=Ne^{-N|c_{n}|^{2}}$ (see Refs.[162, 163]). Quantum typicality [77, 162, 163] implies that the energy expectation value for $|\Psi^{\text{inf}}\rangle$ is exponentially close to the value for the infinite temperature canonical ensemble, which is, in turn, equal to zero [see Eq.(III)]. Since the energy expectation value for $|\Psi^{\text{up}}\rangle$ is also zero, both $|\Psi^{\text{up}}\rangle$ and $|\Psi^{\text{inf}}\rangle$ belong to the same “energy shell” in the sense of the canonical ensemble.
@@
The relaxation plots of $\langle S^{z}_{1}(t)\rangle$ for $S=$ 1/2, 1, 3/2 and 2 are shown in Fig. 16. There one can observe that the thermalization process for $S\geq 3/2$ starting from the initial condition $|\Psi^{\text{up}}\rangle$ is noticeably slower than the one starting from $|\Psi^{\text{inf}}\rangle$. Such a difference is to be expected when atypical quantum scar eigenstates are prominently present in the expansion of $|\Psi^{\text{up}}\rangle$. At the same time, the difference between the two thermalization curves for $S=1/2$ is rather small — suggestive of the absence of quantum scar eigenstates in the expansion of $|\Psi^{\text{up}}\rangle$. The case of $S=1$ appears to be transitional between $S=1/2$ and $S\geq 3/2$.
@@
Figure 17 further illustrates the anomalous slowdown of the thermalization for spin-3/2 chains as a function of chain length $L$. The coincidence of the initial slowed down behavior for the chains of different length indicates that the slowdown of the thermalization process starting from the state $|\Psi^{\text{up}}\rangle$ is a feature that is also present in the thermodynamic limit $L\to\infty$.
@@
We, finally, note that Fig. 16 also includes the results of classical simulations for an ensemble of initial conditions devised to imitate the relaxation of quantum spins $S$ starting from the initial state $|\Psi^{\text{up}}\rangle$. In the simulations, each classical spin has length 1. In order to represent a quantum state $|m=S\rangle$, the range $[-1,1]$ of the $z$-projection of a classical spin is divided into $12S+$ 1 equal intervals of length $\frac{2}{2S+1}$ and then the classical projection $S_{z}$ is randomly sampled from the topmost interval $[1-\frac{2}{2S+1},1]$, while the $x$- and $y$-projections are chosen as $S_{x}=\sqrt{1-S_{z}^{2}}\cos\varphi$ and $S_{y}=\sqrt{1-S_{z}^{2}}\sin\varphi$, where $\varphi$ is a random angle from the interval $[0,2\pi)$. On the one hand, the above ensemble becomes increasingly close to the considered periodic classical trajectories as the value of quantum spin $S$ increases. On the other hand, as one can see in Fig. 16, the average for that ensemble exhibits an excellent agreement with the quantum simulations starting already with $S=1$. This is yet another indication, in addition to the observations of Ref.[68], that classical spin ensembles can quantitatively represent quantum spin dynamics for translationally invariant chains of spins $S\geq 1$ at high temperatures. The classicalization of higher spins was also recently reported at the operator level in [164], further highlighting the difference in dynamical properties between spin-$1/2$ and higher spins. In the present context, the success of the above simulation indicates the proximity of quantum dynamics to the semiclassical limit, which will be put forward below as a condition for the observability of distinct many-body quantum scars.